# Query-Guided Self-Supervised Summarization of Nursing Notes

### v1

@@ 5.5 Discussion
A high-quality summary should facilitate efficient understanding of the relevant content for users. In the context of nursing notes, this means the summary should capture the patient’s condition. Our method employs patient-related queries, indirectly ensuring that the summary centers around the patient’s status. The summaries generated with different queries can be seen as coming from distinct conditional distributions and parts of the semantic space. As discussed in Section 5.1, the queries can guide summaries to focus on specific aspects of the original note. Therefore, by selecting appropriate queries, we can control preferences for desired content and adjust granularity, which facilitates a more flexible and user need -oriented summary generation. For instance, broad queries about the patient’s condition will result in a summary that focuses on the patient’s overall condition, while more detailed queries, such as those regarding cardiovascular health, could produce a summary that focuses on that specific aspect.
@@
One challenge is how to efficiently integrate information into the model while avoiding excessive computational cost. We utilize cross-attention to allow the patient’s metadata to efficiently interact on multiple levels with the process of generating the summary. In contrast, for temporal information in previous notes, using cross-attention in a similar manner might make it difficult for the model to balance attention across the current note, past notes, and patient information, in addition to introducing computational challenges with long sequences of notes. Hence, we adopt a simple but effective strategy: representing the temporal information, obtained by weighted mean pooling from previous note representations, as the first token of the decoder’s input. This strategy is intuitive, as information from previous notes naturally precedes the summary of the current note.
@@
The metrics used in the automatic evaluation have limitations as they do not conclusively reflect the quality of the summary, and come with trade-offs. For example, a good performance in predictiveness and medical information consistency (UMLS-Recall) may not be due to the high quality of the summary but rather caused by copying the source note, resulting in a lack of conciseness and fluency. Conversely, as the summary becomes more concise, it may become less informative. Furthermore, models used to measure factuality and general consistency have inherent biases. As they are based on general semantics, they are potentially weak at recognizing patient-related information due to the dissimilarity between their training domain and clinical data, and they often prioritize text style and structure. Finally, since BARTScore is derived from the BART model, summaries generated by BART have a bias of scoring relatively higher with this metric. We attempt to mitigate the impact of these limitations by comprehensively considering multiple metrics, and including the manual evaluation by a clinician, but there remains a need for more conclusive evaluation metrics.
@@
(1) Our current approach produces summaries of individual nursing notes, lacking the long context and support for multiple note summarization. (2) There is room for more exploration on the formulation of the clinical queries. We don’t employ generative queries but only queries related to the classification of the patient’s status. Also, when investigating the combined effects of multiple queries, further exploration using multi-task learning methods could be beneficial. (3) Due to the workload, the number of notes assessed in the manual evaluation is limited to 25. A larger sample size would allow more conclusive comparisons of the strengths and weaknesses of the methods.
@@ Conclusion.
We presented a novel method for self-supervised nursing note summarization, where the main innovation was the introduction of query guidance, which successfully directed the summaries to include desired content. In the manual evaluation by a professional clinician, our method significantly outperformed a specialized open text summarization model, BART-Large-CNN, in all metrics. Because this model was the base model of our method, the result highlights the usefulness of the novel developments. Of the other baselines, the proprietary GPT-4 had the closest performance to our method and was better than the other baselines. In the automatic evaluation, GPT-4 was better than our method in predictiveness, but, importantly, our method outperformed GPT-4 in factual consistency, having fewer hallucinated facts without sacrificing the correct content. The same trend was seen in the manual evaluation as higher average consistency for our method, although the difference was not statistically significant. Hence, our approach can produce more reliable summaries, clearing obstacles for responsible clinical use of LLMs. From the machine learning perspective, our method demonstrates the feasibility of domain adaptation for pre-trained text summarization models without explicit supervision, and the effectiveness of self-supervised strategies to guide conditional summarization to specific interests.
Appendix A. Additional Details for Implementation
@@ A.2 Data preprocessing
Following prior research (Harutyunyan et al., 2019; Huang et al., 2019), we perform filtering on admission records and nursing notes. Initially, we filter out specific admission cases: (1) cases of in-hospital mortality and admissions categorized as ”NEWBORN”; (2) cases containing diagnosis codes outside HCUP CCS code groups. We retain only admissions containing clinical notes categorized as ”Nursing/other”. Subsequently, we apply a length limit to nursing notes, filtering out those with more than 800 tokens or fewer than 50 tokens. Finally, we filtered out admission cases with more than 100 nursing notes. Nursing notes in these cases typically represent out-of-distribution information or are irrelevant to the care episode.
@@ A.3 Details of Implementation for the Query Responder
Given a nursing note $N$ and its summary $S$ generated by QGSumm:
Contrastive Next Note Prediction.
Two note pairs $(N,N_{pos})$ and $(N,N_{neg})$ are constructed as introduced in Section 4.2.
The learning objective in this case is:
$\operatorname{min}\mathcal{L}_{CrossEntropy}([p_{pos},p_{neg}],[p^{\prime}_{pos},p^{\prime}_{neg}]).$ (12)
Readmission Prediction.
Phenotype Classification.
We obtain the probability distribution of the phenotype from $\operatorname{R}$:
Consequently, the learning objective is:
$\operatorname{min}\mathcal{L}_{CrossEntropy}([p_{1},p_{2},...,p_{25}],[p^{\prime}_{1},p^{\prime}_{2},...,p^{\prime}_{25}]).$ (14)
@@ A.4 Hyperparameter Setting
We present hyperparameter settings of QGSumm and the query responder $\operatorname{R}$ in Table 4. The configuration of hyperparameters for the base model’s architecture keeps the same as the original configuration of BART-Large-CNN [^3] https://huggingface.co/facebook/bart-large-cnn/blob/main/config.json. We set the maximum length of the summary to 500 tokens, allowing for flexibility as we aim for the model to autonomously determine the appropriate length. We use Adam optimizer (Kingma and Ba, 2014) to optimize the model.
@@
Appendix B. Additional Details for Baselines and Evaluation
@@ B.1 Baselines
BART-Large-CNN (Lewis et al., 2020): It is chosen as the base model for its excellent performance on text summarization as well as less computation cost than its peers. We consider it as one baseline in the experiment to illustrate performance without the proposed novel components. Pegasus (Zhang et al., 2020a): It is a transformer-based pre-trained model specialized in abstractive summarization, widely recognized as a baseline model in many studies on text summarization. BioMistral-7B (Labrak et al., 2024): It is an open-source instruction-based LLM adapted from Mistral (Jiang et al., 2023) for the medical domain. It achieves state-of-the-art performance in supervised fine-tuning benchmarks compared to other open-source medical language models. GPT-4 (OpenAI, 2023): It is a proprietary LLM representing state-of-the-art on general NLP task.
Consistent with our method’s settings, the maximum length of the summary is set to 500 tokens.
Prompt Learning for GPT-4 and Bio-Mistral
A prompt example is shown in Figure 7. We employ one-shot in-context learning to prompt GPT-4 and BioMistral-7B, providing one summary example. This strategy can significantly improve the conciseness of the generated summary and ensure its text structure aligns with the requirements.
Figure 7: One prompt example used when testing GPT-4 and BioMistral.
@@
We randomly sample 10 nursing notes from the training set and pair them with their corresponding summaries generated by GPT-4 to create the training data for 10-shot fine-tuning of BART-Large-CNN, BioMistral-7B, and Pegasus. The training data is transformed into instructions-formatted prompts as shown in Figure 7 for fine-tuning BioMistral-7B. We fine-tune BART-Large-CNN and Pegasus for 8 and 9 epochs, respectively, with a learning rate of 0.0005. As for BioMistral-7B, we fine-tune it using QLoRA adaptation for 7 epochs with a learning rate set to 0.0002.
Evaluation on Predictiveness
In our experiments, we assess predictiveness by evaluating the performance of readmission prediction and phenotype classification using summaries from various baselines (including 8 different methods, which also includes extractive methods) and different settings of our approach. For instance, to evaluate the predictiveness of summaries from BART-zs, we begin with the trained query responder $\operatorname{R}$ as described in Appendix A.3 Details of Implementation for the Query Responder. We then fine-tune $\operatorname{R}$ and compute the metrics through 10-fold cross-validation using summaries of nursing notes in the test set generated from BART-zs. Consequently, we obtain multiple responders corresponding to each method for readmission prediction and phenotype classification, respectively.
@@ B.2 Grading Criteria of Manuel Evaluation
Figure 9: Effects of various values of $\lambda_{1}$ (Equation 3) and $\lambda_{2}$ (Equation 7). (a): $\lambda_{1}$ affects both the length and the medical information consistency of the generated summary, resulting in changes in the summary’s predictiveness; (b): $\lambda_{2}$ regulates the importance of the PIA block. We use “Readmission Prediction and Phenotype Classification” as the query. The PIA is applied to all decoder layers.
Appendix C. Additional Experimental Results
Effects of the length penalty coefficient.
@@ Appendix C. Additional Experimental Results
Effects of the importance of patient meta information.

### v2

@@ 5.4 Discussion and Conclusion
A high-quality summary should facilitate efficient understanding of the relevant content by clinical personnel, especially for a nursing note the summary should capture the patient’s condition. Our method employs patient-related queries, indirectly ensuring that the summary centers around the patient’s status. The summaries generated with different queries can be seen as coming from distinct conditional distributions and parts of the semantic space, allowing control over content and granularity. This facilitates a more flexible and user need -oriented summary generation. We show in C.6 that the queries can guide summaries to focus on specific aspects of the original note. For instance, broad queries produce comprehensive summaries, while detailed queries yield focused summaries on specific conditions like cardiovascular health.
@@
One challenge is how to efficiently integrate information into the model without excessive computational cost. We use cross-attention to allow the patient’s metadata to efficiently interact on multiple levels during summary generation. In contrast, for temporal information in previous notes, using cross-attention in a similar manner might make it difficult for the model to balance attention across the current note, past notes, and patient information, and increase computational challenges with long sequences. Hence, we represent the temporal information, obtained by weighted mean pooling from previous note representations, as the first token of the decoder’s input. This strategy is intuitive, as information from previous notes naturally precedes the summary of the current note.
@@
The metrics used in the automatic evaluation have limitations as they do not conclusively reflect the quality of the summary, and come with trade-offs. For example, a good performance in predictiveness and medical information consistency (UMLS-Recall) may not be due to the high quality of the summary but rather caused by copying the source note, resulting in a lack of conciseness and fluency. Conversely, as the summary becomes more concise, it may become less informative. Furthermore, models used to measure factuality have inherent biases. They are potentially weak at recognizing patient-related information due to the dissimilarity between their training domain and clinical data. We attempt to mitigate the impact of these limitations by comprehensively considering multiple metrics, and including the manual evaluation by a clinician. However, there remains a need for more conclusive evaluation metrics, particularly when ground truth is unavailable.
@@
(1) Our current approach produces summaries of individual nursing notes, and lack the long context and support for multiple note summarization. (2) There is room for more exploration on the formulation of the clinical queries, such as generative queries. Also, when investigating the combined effects of multiple queries, further exploration using multi-task learning methods could be beneficial. (3) Due to the workload, the number of summaries assessed in the manual evaluation is limited to 100.
@@ Conclusion.
We presented a novel self-supervised nursing note summarization method, where the main innovation was the introduction of query guidance, which successfully directed the summaries to include desired content. In the manual evaluation by a professional clinician, our method significantly outperformed a specialized open text summarization model, BART-Large-CNN, in all metrics. Of the other baselines, the proprietary GPT-4 had the closest performance to our method and was better than the other baselines. In the automatic evaluation, our method outperformed GPT-4 in factual consistency, having fewer hallucinated facts without sacrificing the correct content. The same trend was seen in the manual evaluation as higher average consistency for our method. Hence, our approach can produce more reliable summaries, clearing obstacles for responsible clinical use of LLMs. Our method demonstrates the feasibility of domain adaptation for pre-trained text summarization models without explicit supervision, and the effectiveness of self-supervised strategies to guide conditional summarization to specific interests.
[…]
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI, and grants 352986, 358246) and EU (H2020 grant 101016775 and NextGenerationEU).
Appendix A Additional Details for Implementation
@@ A.2 Data Preprocessing
Following prior research (Harutyunyan et al., 2019; Huang et al., 2019), we perform filtering on admission records and nursing notes. Initially, we filter out specific admission cases: (1) cases of in-hospital mortality and admissions categorized as ”NEWBORN”; (2) cases containing diagnosis codes outside HCUP CCS code groups. We retain only admissions containing clinical notes categorized as ”Nursing/other”. After the initial filtering, we get the statistics of The distribution of the number of nursing notes per admission, shown in Table LABEL:tab:num_notes. Subsequently, we apply a length limit to nursing notes, filtering out those with more than 800 tokens or fewer than 50 tokens. Finally, we filtered out admission cases with more than 100 nursing notes. Nursing notes in these cases typically represent out-of-distribution information or are irrelevant to the care episode.
@@
[…]
tab:num_notes percentiles number of nursing notes per admission 25% 3 50% 6 75% 13 90% 35 100% 913
Table 3: The distribution of the number of nursing notes per admission.
@@ A.3 Details of Implementation for the Query Responder
In addition to Readmission Prediction and Phenotype Classification, we also employ another query, Contrastive Next Note Prediction. Hence, totally five different queries are used in the experiment.
Contrastive Next Note Prediction.
Given a nursing note pair $(N,N^{\prime})$, we regard the query about whether $N^{\prime}$ is the successor note of $N$ as a prediction of the patient’s future status. To train the query responder R for the next note prediction, we create two note pairs for each nursing note, where the positive pair $(N,N_{pos})$ comprises the note and its successor in the sequence, and the negative pair $(N,N_{neg})$ contains the note and a randomly chosen non-consecutive note. If $N$ is the patient’s last nursing note, we use the patient’s discharge summary and a random note from other patients to construct the positive and negative pairs. The query is formulated as binary classification, and the output of $\operatorname{R}$ is the probability of each pair being the positive pair containing the consecutive notes:
The learning objective in this case is:
$\operatorname{min}\mathcal{L}_{CE}([p_{pos},p_{neg}],[p^{\prime}_{pos},p^{\prime}_{neg}]).$ (7)
Readmission Prediction.
Phenotype Classification.
The prediction process can be formulated as:
Consequently, the learning objective is:
$\operatorname{min}\mathcal{L}_{CE}([p_{1},p_{2},...,p_{25}],[p^{\prime}_{1},p^{\prime}_{2},...,p^{\prime}_{25}]).$ (9)
Readmission Prediction and Phenotype Classification.
Training of $\operatorname{R}$.
@@ A.4 Hyperparameter Setting
We present hyperparameter settings of QGSumm and the query responder $\operatorname{R}$ in Table 5. The configuration of hyperparameters for the base model’s architecture keeps the same as the original configuration of BART-Large-CNN [^3] https://huggingface.co/facebook/bart-large-cnn/blob/main/config.json. We set the maximum length of the summary to 500 tokens, allowing for flexibility as we aim for the model to autonomously determine the appropriate length. We use Adam optimizer (Kingma and Ba, 2014) to optimize the model. The experiments are ran on one Tesla A100 with 80G memory and one Tesla P100 with 16G memory.
@@
A.5 Next Token Prediction
As illustrated in 3.3, we obtain the hidden representation $\mathbf{H}^{dec}$ from the decoder’s final layer. The process of generating the next token then can be abstracted as:
$[\mathbf{H}^{enc},\mathbf{H}^{PA}]=\operatorname{ENC}(N_{i},PA),$ (10)
$\mathbf{H}^{dec}=\operatorname{DEC}(\mathbf{H}^{enc},\mathbf{H}^{PA},[[TI],[BOS],y_{1},\ldots,y_{j}]),$ (11)
$\mathbf{v}=\operatorname{LMH}(\mathbf{H}^{dec}),$ (12)
$\mathbf{v^{\prime}}=\operatorname{ST-GumbelSoftmax}(\mathbf{v}).$ (13)
Appendix B Additional Details for Baselines and Evaluation
@@ B.1 Baselines
BART-Large-CNN (Lewis et al., 2020): It is chosen as the base model for its excellent performance on text summarization as well as less computation cost than its peers. We consider it as one baseline in the experiment to illustrate performance without the proposed novel components. Pegasus (Zhang et al., 2020a): It is a transformer-based pre-trained model specialized in abstractive summarization, widely recognized as a baseline model in many studies on text summarization. BioMistral-7B (Labrak et al., 2024): It is an open-source instruction-based LLM adapted from Mistral (Jiang et al., 2023) for the medical domain. It achieves state-of-the-art performance in supervised fine-tuning benchmarks compared to other open-source medical language models. GPT-4 (OpenAI, 2023): It is a proprietary LLM representing state-of-the-art on general NLP task. We employ one-shot in-context learning to prompt GPT-4 and BioMistral-7B, providing one summary example to ensure the structure of generated summaries aligns with the requirements. We show prompt examples in C.5. Consistent with our method’s settings, the maximum length of the summary is set to 500 tokens.
In Table LABEL:tab:num_parameters, we provide the statistics on the total number of parameters for QGSumm and baselines.
[…]
tab:num_parameters number of parameters QGSumm 457M BART-Large-CNN 406M Pegasus 568M BioMistral 7B
Table 6: The statistics on the total number of parameters for QGSumm and baselines (except GPT-4, whose parameter count remains undisclosed).
We also consider two representative extractive methods, TextRank (Mihalcea and Tarau, 2004) and Lead-40%, and report the evaluation results of them as a reference. We do not compare our methond and baselines with them, since extractive methods are out of the scope of this work. In TextRank, we utilize MPNet (Song et al., 2020) to obtain sentence embeddings. In Lead-40%, we use the first 40% of the content of the note as a summary.
@@
In addition, we show the performance of baselines in few-shot fine-tuning settings for reference. We randomly sample 10 nursing notes from the training set and pair them with their corresponding summaries generated by GPT-4 as the sliver ground truth to create the training data for 10-shot fine-tuning of BART-Large-CNN, BioMistral-7B, and Pegasus. The training data is transformed into instructions-formatted prompts for fine-tuning BioMistral-7B. We fine-tune BART-Large-CNN and Pegasus for 8 and 9 epochs, respectively, with a learning rate of 0.0005. As for BioMistral-7B, we fine-tune it using QLoRA adaptation for 7 epochs with a learning rate set to 0.0002.
@@ B.2 Grading Criteria of Manuel Evaluation
B.3 Evaluations on Predictiveness
Based on the assumption that a high-quality summary, by effectively capturing critical patient information, can better predict both the current and future status of patients, we introduce predictiveness as a novel metric to evaluate the quality of the summaries.
Specifically, we assess predictiveness by measuring the performance of readmission prediction and phenotype classification using summaries generated by various baselines. The evaluation workflow for this metric is as follows:
  • Generate summaries of notes in the test set for each method (baselines and QGSumm);
  • For summaries from one method, split them into 10 folds;
  • Fine-tune the own query responder of each method using its summaries, and compute the metrics through 10-fold cross-validation.
For instance, to evaluate the predictiveness of summaries from BART, we begin with the trained query responder $\operatorname{R}$ as described in A.3. We fine-tune $\operatorname{R}$ using the BART-generated summaries of nursing notes from the test set and then compute the relevant metrics through 10-fold cross-validation. Consequently, we obtain multiple responders tailored to each method, specifically for readmission prediction and phenotype classification.
We fine-tune each method’s query responder using its own summaries to ensure the most accurate evaluation of predictiveness for each method. For comparison, we also experimented with using only GPT-4’s summaries to fine-tune a single query responder across all methods. As shown in Table LABEL:tab:eval_pred, the results got worse for all methods distinct from GPT-4, which might be due to the dissimilarity of the semantic spaces between summaries from GPT-4 and summaries from others. By optimizing a separate predictor for each model we get the best possible predictions for each one, leading to a fair comparison.
[…]
tab:eval_pred Weighted F1 Macro F1 QGSumm $\downarrow$1.6 $\downarrow$0.7 BART $\downarrow$1.9 $\downarrow$0.9 BioMistral $\downarrow$0.9 $\downarrow$0.4 Pegasus $\downarrow$1.7 $\downarrow$0.7
Table 7: Reults on predictiveness get worse when using only summaries from GPT-4 to finetune the query responder. Weighted F1 and Macro F1 are reported for readmission prediction and phenotype classification, respectively.
Figure 7: One artificial nursing note and its summaries from QGSumm, BART, GPT-4, and BioMistral, respectively.
Figure 8: Effects of various values of $\lambda_{1}$ (Equation 3.2) and $\lambda_{2}$ (Equation 3.3). (a): $\lambda_{1}$ affects both the length and the medical information consistency of the generated summary, resulting in changes in the summary’s predictiveness; (b): $\lambda_{2}$ regulates the importance of the PIA block. The PIA is applied to all decoder layers.
Appendix C Additional Experimental Results
C.1 Case Study
One artificial nursing note and its corresponding summaries generated by QGSumm, BART, GPT-4, and BioMistral are presented in Figure 7. In the original nursing note, the content highlighted in blue indicates information included in the summary generated by our approach. We can see that our approach captures most of the important patient information. However, some details, such as cardiovascular and respiratory conditions, are overlooked. The summary from BART only covers information from the first half of the nursing note, suggesting the limitation in understanding long context. Summaries from GPT-4 and BioMistral contain more patient information but lack conciseness. These models achieve fluency by rephrasing notes and expanding abbreviations. However, BioMistral struggles with maintaining factuality, often excessively reasoning about the patient’s personal information and condition.
Table 8: Additional results of: (1) automatic evaluation on few-shot fine-tuning and extractive methods; (2) using five different queries. Lower values are better for Length and UMLS-FDR, higher values for the other metrics. The subscripts denote standard deviation.
Type Method Consistency and Factuality Conciseness Predictiveness
UMLS-Recall UMLS-FDR FactKB Length Readmission Phenotype
Weighted F1 F1 Macro F1
Few-Shot BART 52.57.3 44.57.1 0.760.15 65.0% 82.20.5 14.41.3 21.10.4
BioMistral 57.210.2 49.17.8 0.700.15 68.8% 81.70.4 10.21.1 22.00.4
Pegasus 35.18.4 52.67.7 0.700.18 57.4% 80.50.8 12.51.8 18.30.6
Extractive Lead-40% 42.76.7 0.302.6 0.990.06 40.0% 83.10.6 12.61.5 21.70.5
TextRank 58.57.9 0.081.4 0.950.12 51.9% 81.90.7 14.41.7 23.30.5
QGSumm -Similarity 53.17.2 20.76.7 0.820.13 51.7% 79.50.6 12.01.2 22.40.4
-NextNote 56.48.0 35.27.1 0.770.11 49.3% 80.80.6 11.71.4 23.20.6
-Readmission 58.27.5 22.76.5 0.780.14 46.2% 82.40.5 18.21.6 23.90.5
-Phenotype 58.57.4 36.26.9 0.790.13 48.0% 81.90.5 13.41.5 25.60.6
-Re+Ph 58.87.9 24.16.4 0.800.14 48.2% 84.20.5 17.21.6 25.10.5
C.2 Effects of the Length Penalty Coefficient
@@
C.3 Effects of the Importance of Patient Meta Information
@@
C.4 Results on Few-shot and Extractive Methods
We present additional results on few-shot fine-tuning settings and extractive methods, as shown in Table 8.
After fine-tuning with 10 summaries generated by GPT-4, BART and Pegasus produce summaries that include more medical and patient-related information, as indicated by higher F1-score and UMLS-Recall. However, their performance declines in certain factual metrics and their summaries become less concise, mirroring the performance trend seen with GPT-4. BioMistral’s performance has not changed much after few-shot adaptation. Considering the inherent bias from the limitations of GPT-4, we do not compare QGSumm with few-shot fine-tuning methods in the main text.
Table 9: We conduct evaluations on 100 nursing notes using six different prompts for GPT-4 and BioMistral. Lower values are better for Length and UMLS-FDR, higher values for the other metrics.
UMLS-Recall UMLS-FDR FactKB Length Weighted F1 Macro F1
GPT-4 Original Prompt 56.1 44.9 0.73 56.5% 84.6 23.4
- Patient Information 53.8 48.0 0.76 63.6% 83.2 21.7
- Readmission 55.6 44.4 0.64 54.5% 83.5 24.0
- Penotype 54.4 45.1 0.61 54.9% 83.5 21.1
- Re+Ph 56.1 45.0 0.64 55.7% 84.6 22.6
- Temporal Info 52.2 50.2 0.65 60.3% 82.4 21.5
BioMistral Original Prompt 53.8 50.3 0.69 67.7% 79.7 21.2
- Patient Information 55.1 54.4 0.64 70.1% 78.6 22.0
- Readmission 53.1 50.3 0.63 67.1% 78.2 20.2
- Penotype 52.2 51.2 0.63 65.9% 77.6 20.9
- Re+Ph 53.8 51.7 0.61 66.7% 79.3 21.4
- Temporal Info 50.4 58.6 0.62 68.2% 76.0 19.3
Figure 9: We use five different prompts for evaluating GPT-4 and BioMistral.
C.5 Prompts for GPT-4 and BioMistral
Examples of the prompts used for evaluating GPT-4 and BioMistral are shown in Figure 9. We observe that when a summary example is not included in the prompt, the model sometimes generates outputs that fail to meet the structural requirements, such as producing lists of sentences. This issue is more prevalent when using GPT-4.
To determine the optimal prompt content, we conduct evaluations using six different prompts, with the results shown in Table 9. Given the cost of GPT-4, we limit the evaluation to 100 randomly selected nursing notes from the test set. Compared to the original prompt, adding patient information or queries did not result in significant differences in performance for either GPT-4 or BioMistral. When considering temporal information by including all previous notes in the prompt, we observed that including temporal information interfered with the summarization of the current note, causing both GPT-4 and BioMistral to miss key information from the current note. Consequently, we choose to use the original prompts in the main experiments, as they yielded more balanced results across different metrics.
Furthermore, we note that unlike our method, using prompts that explicitly mention readmission prediction or phenotype classification does not improve performance on those specific tasks. This highlights the effectiveness of using the query to parametrically guide the model’s behavior.
C.6 Effectiveness of the Query Guidance
We provide full results of employing five different queries shown in Table 8. We can observe: (1) Regarding predictiveness, employing queries closely related to patients and focusing on readmission and phenotype information yield superior performance compared to other queries. As expected, the method with phenotype-related queries performs the best in phenotype classification, while the method with readmission-related queries is the best in readmission prediction. This highlights the effectiveness of guiding the summarization with queries, and different queries enable the summary to concentrate on different aspects of the original note. (2) Using similarity as guidance can produce summaries that are more similar to the original notes, resulting in higher scores on general consistency and factuality. However, summaries generated under this configuration tend to be longer and often sacrifice predictiveness and informativeness regarding medical concepts, demonstrating the limitations of the unconstrained guidance signal. (3) Using Next Note Prediction as the query results in weaker performance compared to patient-related queries, indicating that queries lacking meaningful context fail to guide the model effectively. (4) When employing joint guidance with both readmission and phenotype information, our method consistently achieves excellent performance across all metrics. This indicates that combining different guidance signals can help in producing better summaries, and further research is needed to explore this aspect in depth.